OpenStudy (anonymous):

how do I solve 2x-8y=15, -6x+24y=15

OpenStudy (anonymous):

system of equations

OpenStudy (uri):

2x-8y=15 (mutiply by 3) -6x+24y=15 6x-24y=45 -6x+24y=15 ========== -6 +6 will be cancelled -24 +24 will be cancelled No solution

OpenStudy (anonymous):

So I multiply just the two by 3?

OpenStudy (uri):

Noo the whole first eq

OpenStudy (anonymous):

okay so I write it like this 2x*3-8y*3=15 6x-24y=15 but when i subtracted the 6 from the 24 i came up with 18

OpenStudy (uri):

we multiply 15 with 3 aswell

OpenStudy (uri):

and why are you subracting 6 from 24...we add first eq and second eq

OpenStudy (uri):

this has no solution anyway